5 3/4 = d - 2 1/8
3/4 = .75
5.75
8/1 = .125
2.125
-5.75- 2.125 = -3.625
but dont know where to go from there

Answer by Alan3354(69443)

5 3/4 = d - 2 1/8
-------------
3/4 = .75
5.75
8/1 = .125 **** you mean 1/8
2.125
-5.75- 2.125 = -3.625
-----
You left out the d
5.75 = d - 2.125
Add 2.125 to both sides
d = 5.75 + 2.125
d = 7.875
-----------
or
5 3/4 = d - 2 1/8
23/4 = d - 17/8
Multiply thru by 8
46 = 8d - 17
63 = 8d
d = 63/8